10. Hypothesis Test for Two Populations including t-Test for μ1-μ1 t-Test for μd F-Test for We...

10.

Hypothesis Test for Two Populations including

  • t-Test for μ11
  • t-Test for μd
  • F-Test for

We are interested in determining whether or not the variances of the sales at two small grocery stores are equal. A sample of 21 days of sales at Store A and a sample of 16 days of sales at Store B indicated the following

Store A Store B
nA=21 nB=16
SA=28.284

SB=20

Which of the following is critical values of F at 95% confidence?

Group of answer choices

A. a and d

B. 2.57

C. 0.3891

D. 0.3623

E. 2.76
Answer #1

Answer: B) 2.57

Explanation:

The degrees of freedom for numerator = 21-1 =20

The degrees of freedom for denominator =16-1 =15

The critical values of F at 95% confidence and (20, 15) df =2.57